OpenStudy (anonymous):

If the parabola given by the equation x2 = 4(y − 2) is shifted two units to the right, what will the translated equation be? (x + 2)2 = 4(y − 2) (x + 2)2 = 4(y − 4) (2x)2 = 4(y − 2) (x − 2)2 = 4(y − 2)

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@rational

OpenStudy (michele_laino):

please wait

OpenStudy (campbell_st):

can you tell me the vertex of the parabola..?

OpenStudy (anonymous):

doesn't give me the vertex

OpenStudy (campbell_st):

well the equation is in vertex form \[(x - h)^2 = 4a(y - k)\] (h, k) is the vertex. so any thoughts on the vertex..?

OpenStudy (campbell_st):

p.s. a is the focal length, so in your question a = 1 but you don't need to know it..

OpenStudy (anonymous):

2?

OpenStudy (campbell_st):

is that h or k... and you need 2 values for the ordered pair in the vertex

OpenStudy (anonymous):

(1,2)

OpenStudy (campbell_st):

if that was the case the equation would be \[(x -1)^2 = 4(y - 2)\] I think you need to look at it again

OpenStudy (anonymous):

so the answer is A

OpenStudy (campbell_st):

if that's what you think then select it.... I'm just trying to get you to understand how things are related, and not provide you with answers

OpenStudy (anonymous):

I had got a

OpenStudy (michele_laino):

I think that the equation of traslation, are: \[\left\{ \begin{gathered} x = X - 2 \hfill \\ y = Y \hfill \\ \end{gathered} \right.\] where X, and Y are the new coordinates, or XOY is the new system of coordinates

OpenStudy (campbell_st):

this is really a 3 step process to solving the question 1. identify the current vertex, 2. plot the new vertex, 3. substitute the values into the standard form of the equation to get an answer

OpenStudy (anonymous):

so ima try to find the vertex again

OpenStudy (campbell_st):

that's the 1st step, the vertex of the current equation

OpenStudy (campbell_st):

@Michele_Laino I thought giving answers was contrary to Open Study policies... or is that a privilege of of being a "Qualified Helper"?

OpenStudy (anonymous):

So the answer is A

OpenStudy (campbell_st):

If that's you choice... select it I'm just trying to help improve your understanding...

OpenStudy (anonymous):

I got a I hope I dont get the wrong answer

OpenStudy (campbell_st):

the equation you are given is already in vertex form which is \[(x - h)^2 = 4a(y - k)\] you have \[x^2 = 4(y - 2)\] so you where correct earlier k = 2 what is the value of h..?

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

no... the standard for is \[(x - h)^2~~~~you ~only ~have ~~x^2~~~~ what ~could ~h ~be?\]

OpenStudy (michele_laino):

Please note that I haven't gave the direct answer, and I have no any privilege @campbell_st

OpenStudy (anonymous):

so h is 1

OpenStudy (campbell_st):

lol...@Michele_Laino so this isn't an answer..? I think that the equation of traslation, are: {x=X−2y=Y where X, and Y are the new coordinates, or XOY is the new system of coordinates

OpenStudy (anonymous):

So the answer is D

OpenStudy (anonymous):

im kmsl

OpenStudy (campbell_st):

if h = 1 it you be \[(x - h)^2~~~and~~~(x - 1)^2\] but you want \[x^2\] any thoughts

OpenStudy (campbell_st):

if that's what you think, then select it

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

im 100% sure its D

OpenStudy (campbell_st):

great then select it... you don't need me to give you the ok

OpenStudy (campbell_st):

but my question is... why did you select D

OpenStudy (anonymous):

I do need you lol

OpenStudy (anonymous):

and it is D